OpenStudy (dragon123mama):

use the quadratic formula to solve the equation x^2 =7x - 6 = 0 I don't remember how to solve something like this. please help.

OpenStudy (mariamikayla):

it's x^2 +7x-6=0?

OpenStudy (dragon123mama):

yeah

OpenStudy (dragon123mama):

no -7x sorry type

OpenStudy (dragon123mama):

typo**

OpenStudy (mariamikayla):

x^2-7x-6=0 a=1 b= -7 c= -6 Delta= b^2 -4ac Delta= (-7)^2 -4*1*(-6) = 49 +24= 73 x1;x2= (-b +/- root from Delta)/2a x1= (-b+ root from Delta)/2a= (-(-7) + root 73)/2*1= (7+ root 73)/2 x2=(-b - root from Delta)/2a= (-(-7) - root 73)/2*1= (7 - root 73)/2 Solution is S={ (7-root 73)/2 ; (7+root 73)/2}

OpenStudy (dragon123mama):

what is root and delta. i have no idea what that is.... you made it more confusing..

OpenStudy (mariamikayla):

\[\sqrt{x}\] -is root

OpenStudy (dragon123mama):

oh

OpenStudy (mariamikayla):

\[\Delta \] -is delta

OpenStudy (dragon123mama):

ok. i had tagged you in another question i needed help with. but you never looked at that one.

OpenStudy (mariamikayla):

i'm still getting used with this site, i'm just a day old :))) i didn't know where to look :))

OpenStudy (dragon123mama):

the hammer at the upper left is notificatins.

OpenStudy (dragon123mama):

Will award medal. i don't know how to solve this. What is the vertex form of the equation? y= -X^2 +12x -4

OpenStudy (dragon123mama):

this is the question.

OpenStudy (mariamikayla):

i've seen that,but i don't know how to see where you've mentioned me :))

OpenStudy (dragon123mama):

i tagged you in it. there would be a notification for that.

OpenStudy (mariamikayla):

yes,it is,but how can i see the question where i'm tagged in? :))

OpenStudy (dragon123mama):

you click on it silly. and it takes you to the question.

OpenStudy (mariamikayla):

aa :))))